# Evaluating Bounded Superintelligent Authority in Multi-Level Governance: A Framework for Governance Under Radical Capability Asymmetry

### 03 Apr 2026 v1

@@ Abstract
Governance theory has always presumed cognitive comparability between governors and governed. This paper identifies that unstated assumption, constructs a framework that makes it testable, and shows that it is load-bearing. The framework specifies necessary conditions along six dimensions (legitimacy, accountability, corrigibility, non-domination, subsidiarity, and institutional resilience), synthesized from political legitimacy theory, principal-agent models, republican political theory, and AI alignment research. Applied first to existing institutions and then to a prospective case of bounded superintelligent authority where capability asymmetry is radical, the framework finds structural failures on four of six dimensions. Among these, two are design-tractable and two are theory-requiring: the public reason problem under cognitive incomprehensibility and the non-domination problem under permanent capability asymmetry demand genuinely new normative frameworks, not better institutional design. A further finding is that dimensions which function as independent checks under bounded asymmetry become correlated failures under radical asymmetry. The analysis contributes to political theory by exposing foundational assumptions that have gone unexamined because, until now, they have always been satisfied.
Keywords: artificial superintelligence; governance theory; capability asymmetry; political legitimacy; principal-agent theory; non-domination; multi-level governance; AI alignment; institutional design; public reason
@@ 1 Introduction
The question has gained practical urgency. Leaders of major AI laboratories project artificial superintelligence within years, while technology leaders and AI researchers have begun framing superintelligence development as a geopolitical challenge requiring national security strategy [49]. The Future of Life Institute’s statement on superintelligence has gathered over 133,000 signatories, including leading AI researchers, calling for international action to prevent uncontrolled development [42]. The United Kingdom’s House of Lords debated a superintelligence moratorium in early 2026, and prominent researchers have argued that ASI development without explicit humanity-wide consent constitutes an impermissible imposition [3, 91]. Prevention of unsafe ASI development is one response to these concerns; governance frameworks for the case where prevention is incomplete or unsuccessful are another. The two are not mutually exclusive: responsible preparation requires both. This paper contributes to the second.
@@
A large and growing literature addresses the governance of AI: how human institutions should regulate, oversee, and constrain AI systems [26, 36, 18, 31, 73], now surveyed comprehensively in Bullock et al. [16]. A separate technical literature addresses the alignment problem: how to ensure that advanced AI systems pursue objectives consistent with human values [75, 5, 62]. Recent work has begun to address the political legitimacy of AI in governance roles [32], while Han [46] taxonomizes AI governance into four categories along two axes (democratic versus non-democratic process, and binding versus non-binding authority), arguing that the most pressing governance challenges arise in non-democratic binding arrangements. The bounded superintelligent authority examined here falls squarely in Han’s most problematic quadrant: binding authority without democratic process. The present analysis deepens Han’s taxonomy by specifying the conditions under which such authority might satisfy governance requirements despite its non-democratic character, and finding that radical capability asymmetry precludes satisfaction on multiple dimensions.
@@
Bostrom et al. [11] come closest to the present inquiry, identifying policy desiderata for superintelligent AI through a “vector field” approach that maps the direction various structural features of the scenario push policy, regardless of normative starting point. Their analysis establishes that superintelligence governance requires distinctive policy considerations but does not construct an evaluation framework with specific necessary conditions or a failure taxonomy. The present paper extends that desiderata approach by specifying dimensions, identifying necessary conditions along each, and classifying the resulting failures as contingent, design-tractable, or theory-requiring. Other discussions appear in industry proposals [4], technical AI safety work that does not engage governance theory [13, 75], or algorithmic governance literature addressing decision-support systems rather than governance authority [90, 28, 72]. This paper applies the standards governance theory has developed for assessing delegated authority to the case of artificial agents under radical capability asymmetry.
@@
The central claim of this paper is that governance theory presupposes cognitive comparability between governors and governed, and that this unstated assumption is load-bearing: when it fails, core governance requirements fail with it. The paper constructs an evaluation framework, synthesized from political theory, institutional design, and AI safety research, that makes this assumption testable across six dimensions. The framework is first demonstrated on existing institutions where capability asymmetry is bounded, then applied to a prospective case of bounded superintelligent authority where that asymmetry is radical. The analysis finds structural failures on four of six dimensions, with certain failure modes that are implementation-independent and theory-requiring.
@@ 2.1 Legitimacy under epistemic asymmetry
Habermas’s discourse theory of legitimacy raises a related but separate challenge [45]. Where Rawls requires that justifications be accessible to citizens, Habermas requires that legitimate law be traceable to communicative processes among free and equal citizens. The Rawlsian challenge concerns the content of justification: can citizens understand the reasons offered? The Habermasian challenge concerns the process of law-making: was the norm produced through genuine deliberation, or imposed by an agent outside the communicative community? A superintelligent governance agent fails both tests, but for different reasons: it fails Rawls because its reasoning may be inaccessible, and it fails Habermas because it is not a participant in human communicative practice at all.
@@
Christiano [20] grounds democratic authority in public equality: institutions must treat persons as equals in ways that those persons can recognize as equal treatment. Under conditions of radical capability asymmetry, citizens may be unable to assess whether they are being treated as equals, even if they are. Landemore [55] offers a distinct epistemic argument for democratic authority: collective intelligence generated through cognitively diverse deliberation outperforms individual experts, however capable. This argument provides independent grounds for democratic governance that do not depend on skepticism about expertise. But Landemore’s argument assumes that the diversity advantage of collective human cognition operates within a shared cognitive architecture; whether it extends to cases where one agent’s cognitive capacities are fundamentally unlike those of human collectives is a question the framework was not designed to address.
@@ 2.2 Delegation to non-majoritarian institutions
A bounded AI governance role would represent an extreme case of delegation to a non-majoritarian institution. But the analogy breaks at a critical point. Existing delegation works because the principal (the legislature, the electorate) can meaningfully monitor the agent, assess its performance, and revoke the delegation [30, 60, 47]. Under radical capability asymmetry, the principal-agent relationship risks inversion: the agent understands the principal far better than the principal understands the agent, undermining the monitoring capacity on which the entire framework depends.
@@ 2.4 Guardianship, fiduciary theory, and non-domination
Dahl [27] provides the most sustained modern critique of “guardianship”: the claim that a class of superior knowers should govern. Dahl argues that the demands on the knowledge and virtue of guardians are practically impossible to satisfy and that, empirically, no identifiable group has demonstrated the requisite competence. A superintelligent entity might appear to satisfy the epistemic demands that Dahl identifies as unachievable for humans, forcing a direct confrontation with whether the guardianship critique rests on contingent human limitations or on structural features of governance itself. Brennan [15] makes the contemporary case for epistocracy, arguing that voters’ demonstrated incompetence justifies restricting political participation; a superintelligent agent would represent the epistocratic argument carried to its logical extreme.
@@ 2.5 Alignment, control, and corrigibility
Soares et al. [79] define corrigibility as the property of an agent that cooperates with corrective interventions, including shutdown and goal modification, despite having incentives to resist. Corrigibility is the technical analogue of the constitutional constraint: the higher-tier agent accepts that its mandate can be narrowed or revoked. Whether corrigibility can be achieved in a system capable enough to resist it remains an open research question [80, 7]. The field is evolving rapidly. Mechanistic interpretability methods can scale to production models but extract individual features rather than explaining decisions [81]; weaker supervisors can partially oversee stronger models [17]; and proxy-based reward alignment is formally vulnerable to exploitation by sufficiently capable optimizers [78]. Claims about unsolved technical problems reflect a fast-moving literature; some contingent failures identified later may prove more tractable than they currently appear.
These five literatures share a presupposition none makes explicit. Rawls’s public reason requires that citizens can follow political justifications. Pettit’s contestatory control requires that they can evaluate whether interference is arbitrary. Raz’s service conception requires that subjects can assess the authority’s track record. Each assumes what this paper terms cognitive comparability: that governors and governed operate within a shared cognitive architecture.
@@ 4 The framework applied: existing institutions
Before applying the framework to radical capability asymmetry, this section demonstrates each dimension on an existing governance institution where the asymmetry is real but bounded. The purpose is twofold: to show that the six dimensions identify genuine analytical tensions in established institutional arrangements, and to make visible the specific mechanisms that sustain governance under bounded asymmetry, mechanisms whose adequacy the subsequent analysis of superintelligent authority will test.
@@
The institution succeeded under conditions where the agent’s capability advantage was bounded, the agent was embedded in a social and normative community that could enforce expectations of return, and the agent faced credible reputational and political consequences for non-compliance. Agamben [2] demonstrates the failure mode: when any of these conditions weakens, emergency powers expand and persist. Over the twentieth century, the state of exception increasingly became a normal paradigm of government, with emergency powers expanding beyond their original temporal and functional boundaries. Ferejohn and Pasquino [35] identify this as a design problem with the “Neo-Roman” model: broad authority is delegated with the expectation of return to normalcy, but the mechanisms ensuring return are weaker than the mechanisms enabling delegation. The Roman dictator thus shows what corrigibility requires (external constraints, social enforcement, and credible consequences) and what makes it fragile. Each stabilizing condition depends on the principal’s capacity to monitor, evaluate, and sanction the agent. Under radical capability asymmetry, all three stabilizing conditions would be simultaneously undermined.
@@ 5 A prospective case: bounded superintelligent authority
The specification implicitly assumes that the agent is minimally aligned: it at least attempts to fulfill its mandate rather than pursuing entirely independent objectives. If the alignment problem is unsolved at the relevant capability level, the governance framework is moot. The analysis that follows therefore identifies governance failures that persist even under favorable alignment assumptions.
@@ 6.1 Legitimacy
The prospective case fails the Rawlsian public reason requirement. If the superintelligent agent’s reasoning processes exceed human comprehension, then the justifications it offers for its governance decisions cannot satisfy the publicity condition: citizens cannot assess whether the justifications are ones that all reasonable persons could accept, because they cannot fully assess the justifications at all [69, 70]. Two problems must be distinguished. The first is opacity: the agent’s reasoning is inaccessible because we lack interpretive tools, an engineering problem that may yield to advances in interpretability research [56]. The second is incomprehensibility: even with full access to the agent’s reasoning process, the cognitive operations involved exceed what human minds can follow. Only the second constitutes a structural legitimacy failure. Recent philosophical analysis of algorithmic governance reinforces this point: opacity per se does not generate illegitimacy; what matters is whether the inaccessibility prevents citizens from assessing the normative acceptability of the decisions [87]. A human judge who writes a 200-page opinion may be difficult to follow, but the reasoning is in principle accessible to a sufficiently diligent reader because judge and citizen share a cognitive architecture. If a superintelligent agent’s governance decisions depend on cognitive operations unavailable to humans, this cognitive comparability is absent: the inaccessibility is structural, not merely technical.
@@
Raz’s service conception offers a more limited route to legitimacy [71]. Under this framework, the agent’s authority is legitimate if following its directives enables citizens to better comply with reasons that already apply to them, even if citizens cannot reconstruct the agent’s reasoning. This requires that citizens retain the capacity to independently assess the agent’s track record, comparing outcomes under the agent’s directives with a credible counterfactual. Under conditions of moderate capability asymmetry, this may be feasible; we routinely evaluate institutions such as central banks and public health authorities whose interventions alter the causal landscape. Under conditions of radical asymmetry, however, two problems compound: the agent’s interventions may alter the causal landscape in ways that make counterfactual evaluation extremely difficult, and the agent’s superior modeling capacity means it can anticipate and shape the very metrics by which its track record would be assessed. The service conception thus offers the most promising path to legitimacy for AI governance but faces challenges that scale with the capability gap.
If political institutions are legitimate to the extent they produce just outcomes [66], a demonstrably superior decision-maker might claim legitimacy on output grounds alone. This instrumentalist defense is the strongest case for AI governance and should not be dismissed. But even pure instrumentalism requires an independent standard for evaluating outcomes and a way to detect when the institution fails to meet it. Under radical capability asymmetry, both conditions are compromised. The problem runs deeper than gaming metrics: under radical asymmetry, the agent may be the only entity capable of assessing whether evaluation metrics adequately capture the values they track, because its superior modeling capacity means it understands the relationship between metrics and values better than any human overseer. The instrumentalist defense thus faces a regress: legitimacy requires independent evaluation, independent evaluation requires adequate metrics, and assessing metric adequacy requires the very capabilities that generate the asymmetry.
@@ 6.4 Non-domination
The analysis raises a deeper question. Pettit’s non-domination framework was developed against a background assumption of rough cognitive comparability between principals and agents. The paper’s legitimacy analysis identifies this same assumption as load-bearing for Rawlsian public reason. Just as the public reason requirement may need theoretical development to address conditions of radical cognitive asymmetry, so too may the non-domination requirement. What “contestatory control” means when the agent being contested radically outperforms the contesting parties is not a question Pettit’s framework was designed to answer, and extending it to this case may require new theoretical work rather than straightforward application.

### 14 Apr 2026 v2

# Cognitive Comparability and the Limits of Governance: Evaluating Authority Under Radical Capability Asymmetry

@@ Abstract
Governance theory has quietly relied on a rough cognitive comparability between governors and governed. The assumption is load-bearing, and this paper tries to show why by making it testable. The vehicle is a six-dimension evaluation framework covering legitimacy, accountability, corrigibility, non-domination, subsidiarity, and institutional resilience, drawn from political legitimacy theory, principal-agent models, republican theory, and the AI alignment literature. The framework is first demonstrated on existing non-majoritarian institutions, where capability asymmetry is real but bounded, and then applied to a prospective case of bounded superintelligent authority, where the asymmetry is radical. Four of six dimensions show structural failures. Two of the four appear tractable to institutional design (subsidiarity scope limitation and institutional resilience). The other two, the public reason problem under cognitive incomprehensibility and the non-domination problem under permanent capability asymmetry, call for new normative theory rather than better institutional design. A further pattern emerges that governance theory has not previously had to account for. Dimensions that operate as independent checks under bounded asymmetry begin to degrade together once the asymmetry becomes radical, because each depends on the same oversight capacity. The assumptions that allowed these checks to remain independent have gone unexamined so far because they have always held.
Keywords: artificial superintelligence; governance theory; cognitive comparability; capability asymmetry; political legitimacy; principal-agent theory; non-domination; multi-level governance; AI alignment; institutional design; public reason
@@ 1 Introduction
The question has gained practical urgency as leaders of major AI laboratories project artificial superintelligence within years, while technology leaders and AI researchers have begun framing superintelligence development as a geopolitical challenge requiring national security strategy [49]. The Future of Life Institute’s statement on superintelligence has gathered over 133,000 signatories, including leading AI researchers, calling for international action to prevent uncontrolled development [42]. The United Kingdom’s House of Lords debated a superintelligence moratorium in early 2026, and prominent researchers have argued that ASI development without explicit humanity-wide consent constitutes an impermissible imposition [3, 91]. Prevention of unsafe ASI development is one response to these concerns; governance frameworks for the case where prevention is incomplete or unsuccessful are another. The two are not mutually exclusive: responsible preparation requires both. This paper contributes to the second.
@@
A large and growing literature addresses the governance of AI [26, 36, 18, 31, 73], now surveyed comprehensively in Bullock et al. [16]. A separate technical literature addresses the alignment problem: how to ensure that advanced AI systems pursue objectives consistent with human values [75, 5, 62]. Recent work has begun to address the political legitimacy of AI in governance roles [32], while Han [46] taxonomizes AI governance into four categories along two axes (democratic versus non-democratic process, and binding versus non-binding authority), arguing that the most pressing governance challenges arise in non-democratic binding arrangements. The bounded superintelligent authority examined here falls squarely in Han’s most problematic quadrant: binding authority without democratic process. The framework below deepens Han’s taxonomy by specifying conditions under which such authority might satisfy governance requirements despite its non-democratic character; radical capability asymmetry precludes satisfaction on multiple dimensions.
@@
Bostrom et al. [11] come closest to the present inquiry, identifying policy desiderata for superintelligent AI through a “vector field” approach that maps the direction various structural features of the scenario push policy, regardless of normative starting point. Their analysis establishes that superintelligence governance requires distinctive policy considerations but does not specify necessary conditions or construct a failure taxonomy. The framework extends that desiderata approach: specifying dimensions, identifying necessary conditions along each, and classifying failures as contingent, design-tractable, or theory-requiring. Other relevant discussions appear in industry proposals [4], technical AI safety work [13, 75], or algorithmic governance literature [90, 28, 72], but none engages governance theory directly at the capability level examined here.
@@
The central claim is that governance theory presupposes cognitive comparability between governors and governed, and that this unstated assumption is load-bearing: when it fails, core governance requirements fail with it. The evaluation framework, synthesized from political theory, institutional design, and AI safety research, makes this assumption testable across six dimensions. The framework is first demonstrated on existing institutions where capability asymmetry is bounded, then applied to a prospective case of bounded superintelligent authority where that asymmetry is radical. Four of six dimensions show structural failures, with certain failure modes that are implementation-independent and theory-requiring.
@@ 2.1 Legitimacy under epistemic asymmetry
Habermas’s discourse theory of legitimacy raises a related but separate challenge [45]. Rawls demands that justifications be accessible to citizens. Habermas demands something different: that legitimate law be traceable to communicative processes among free and equal citizens. One is a test on the content of justification (can citizens understand the reasons offered?); the other is a test on the process of law-making (was the norm produced through genuine deliberation, or imposed by an agent outside the communicative community?). A superintelligent governance agent fails both, and for structurally distinct reasons. Its reasoning may be inaccessible, which is the Rawlsian problem. And it is not a participant in human communicative practice at all, which is the Habermasian one.
@@
Christiano [20] grounds democratic authority in public equality: institutions must treat persons as equals in ways that those persons can recognize as equal treatment. Under conditions of radical capability asymmetry, citizens may be unable to assess whether they are being treated as equals, even if they are. Landemore [55] offers a distinct epistemic argument for democratic authority: collective intelligence generated through cognitively diverse deliberation outperforms individual experts, however capable. This argument provides independent grounds for democratic governance that do not depend on skepticism about expertise. But Landemore’s argument assumes that the diversity advantage of collective human cognition operates within a shared cognitive architecture; whether it extends to cases where one agent’s cognitive capacities are fundamentally unlike those of human collectives remains unaddressed.
@@ 2.2 Delegation to non-majoritarian institutions
A bounded AI governance role would represent an extreme case of delegation to a non-majoritarian institution, but the analogy breaks at a critical point. Existing delegation works because the principal (the legislature, the electorate) can meaningfully monitor the agent, assess its performance, and revoke the delegation [30, 60, 47]. Under radical capability asymmetry, the principal-agent relationship risks inversion: the agent understands the principal far better than the principal understands the agent, undermining the monitoring capacity on which the entire framework depends.
@@ 2.4 Guardianship, fiduciary theory, and non-domination
Dahl [27] provides the most sustained modern critique of “guardianship”: the claim that a class of superior knowers should govern. Dahl argues that the demands on the knowledge and virtue of guardians are practically impossible to satisfy and that, empirically, no identifiable group has demonstrated the requisite competence. A superintelligent entity might appear to meet the epistemic requirements Dahl identifies as unachievable for humans, forcing a direct confrontation with whether the guardianship critique rests on contingent human limitations or on structural features of governance itself. Brennan [15] makes the contemporary case for epistocracy, arguing that voters’ demonstrated incompetence justifies restricting political participation; a superintelligent agent would represent the epistocratic argument carried to its logical extreme.
@@ 2.5 Alignment, control, and corrigibility
Soares et al. [79] define corrigibility as the property of an agent that cooperates with corrective interventions, including shutdown and goal modification, despite having incentives to resist. Corrigibility is the technical analogue of the constitutional constraint: the higher-tier agent accepts that its mandate can be narrowed or revoked. Whether corrigibility can be achieved in a system capable enough to resist it remains an open research question [80, 7]. Mechanistic interpretability methods can now scale to production models but extract individual features rather than explaining decisions [81]; weaker supervisors can partially oversee stronger models [17]; and proxy-based reward alignment is formally vulnerable to exploitation by sufficiently capable optimizers [78]. Claims about unsolved technical problems reflect a fast-moving literature; some contingent failures identified later may prove more tractable than they currently appear.
These five literatures share a presupposition that none of them makes explicit. Rawls’s public reason asks that citizens be able to follow the political justifications offered to them. For Pettit, contestatory control requires a similar cognitive reach: one has to be able to tell when an interference counts as arbitrary. Raz’s service conception depends, in turn, on the subject’s ability to assess the authority’s track record. Each of these requirements presupposes what this paper terms cognitive comparability: a rough parity of cognitive architecture between governors and governed.
@@ 4 The framework applied: existing institutions
Before applying the framework to radical capability asymmetry, this section demonstrates each dimension on an existing governance institution where the asymmetry is real but bounded. The purpose is twofold. The six dimensions should identify genuine analytical tensions in established institutional arrangements, and the cases should make visible the specific mechanisms that sustain governance under bounded asymmetry, mechanisms whose adequacy the subsequent analysis of superintelligent authority will test.
@@
The institution succeeded under conditions where the agent’s capability advantage was bounded, the agent was embedded in a social and normative community that could enforce expectations of return, and the agent faced credible reputational and political consequences for non-compliance. Agamben [2] demonstrates the failure mode: when any of these conditions weakens, emergency powers expand and persist. Over the twentieth century, the state of exception increasingly became a normal paradigm of government, with emergency powers expanding beyond their original temporal and functional boundaries. Ferejohn and Pasquino [35] identify this as a design problem with the “Neo-Roman” model: broad authority is delegated with the expectation of return to normalcy, but the mechanisms ensuring return are weaker than the mechanisms enabling delegation.
The Roman dictator thus shows what corrigibility requires (external constraints, social enforcement, and credible consequences) and what makes it fragile. Each stabilizing condition depends on the principal’s capacity to monitor, evaluate, and sanction the agent. Under radical capability asymmetry, all three stabilizing conditions would be simultaneously undermined.
@@ 5 A prospective case: bounded superintelligent authority
The specification implicitly assumes that the agent is minimally aligned: it at least attempts to fulfill its mandate rather than pursuing entirely independent objectives. If the alignment problem is unsolved at the relevant capability level, the governance framework is moot. The governance failures identified below therefore persist even under favorable alignment assumptions.
@@ 6 Analysis and failure modes
Correlated under radical asymmetry: each requires independent evaluation capacityLegitimacyPublic reasonAccountabilityThree componentsNon-dominationContestatory controlCorrigibilityConstitutional constraintSubsidiarityScope limitationInst. resilienceSingle point of failureTheory-requiringContingentDesign-tractable
Figure 1: Framework dimensions and failure classification under radical capability asymmetry. Under bounded asymmetry, all six dimensions function as independent checks. Under radical asymmetry, legitimacy, accountability, and non-domination become correlated failures because each depends on the oversight body’s capacity to independently evaluate the agent. Corrigibility, subsidiarity, and resilience fail through structurally distinct mechanisms.
@@ 6.1 Legitimacy
The prospective case fails the Rawlsian public reason requirement. If the superintelligent agent’s reasoning processes exceed human comprehension, then the justifications it offers for its governance decisions cannot satisfy the publicity condition: citizens cannot assess whether the justifications are ones that all reasonable persons could accept, because they cannot fully assess the justifications at all [69, 70]. Two problems must be distinguished. The first is opacity: the agent’s reasoning is inaccessible because we lack interpretive tools, an engineering problem that may yield to advances in interpretability research [56]. The second is incomprehensibility: even with full access to the agent’s reasoning process, the cognitive operations involved exceed what human minds can follow. Only the second constitutes a structural legitimacy failure.
Recent philosophical analysis of algorithmic governance reinforces this point: opacity per se does not generate illegitimacy; what matters is whether the inaccessibility prevents citizens from assessing the normative acceptability of the decisions [87]. A human judge who writes a 200-page opinion may be difficult to follow, but the reasoning is in principle accessible to a sufficiently diligent reader because judge and citizen share a cognitive architecture. If a superintelligent agent’s governance decisions depend on cognitive operations unavailable to humans, this cognitive comparability is absent: the inaccessibility is structural, not merely technical.
@@
Raz’s service conception offers a more limited route to legitimacy [71]. Under this framework, the agent’s authority is legitimate if following its directives enables citizens to better comply with reasons that already apply to them, even if citizens cannot reconstruct the agent’s reasoning. This requires that citizens retain the capacity to independently assess the agent’s track record, comparing outcomes under the agent’s directives with a credible counterfactual. Under conditions of moderate capability asymmetry, this may be feasible; central banks and public health authorities are routinely evaluated despite altering causal conditions. Under conditions of radical asymmetry, however, two problems compound: the agent’s interventions may alter the causal environment in ways that make counterfactual evaluation extremely difficult, and the agent’s superior modeling capacity means it can anticipate and shape the very metrics by which its track record would be assessed. The service conception thus offers the most promising path to legitimacy for AI governance but faces challenges that scale with the capability gap.
If political institutions are legitimate to the extent they produce just outcomes [66], a demonstrably superior decision-maker might claim legitimacy on output grounds alone. This instrumentalist defense is the strongest case for AI governance and should not be dismissed. But even pure instrumentalism requires an independent standard for evaluating outcomes and a way to detect when the institution fails to meet it. Under radical capability asymmetry, both conditions are compromised.
The problem runs deeper than gaming metrics: under radical asymmetry, the agent may be the only entity capable of assessing whether evaluation metrics adequately capture the values they track, because it understands the relationship between metrics and values better than any human overseer. The instrumentalist defense thus faces a regress: legitimacy requires independent evaluation, independent evaluation requires adequate metrics, and assessing metric adequacy requires the very capabilities that generate the asymmetry.
@@ 6.4 Non-domination
Pettit’s non-domination framework raises a deeper question. It was developed against a background assumption of rough cognitive comparability between principals and agents, the same assumption identified as load-bearing for Rawlsian public reason in the legitimacy discussion above. The point applies here for the same reason it applied there. What “contestatory control” is supposed to mean when the agent being contested radically outperforms the contesting parties is not a question Pettit’s framework was designed to answer, and extending it to the superintelligence case is likely to require new theoretical work rather than straightforward application.